# Recent Advances in Discrete Speech Tokens: A Review

### v2

@@ VI-D3 Limitations
VII Challenges and Future Directions
@@
VII-1 Low-Bitrate Tokens
For bitrates of tokens, factors $Q$ (number of quantizers) and $F$ (frame rate) play a more important role than $V$ (vocabulary size). Using only a single codebook is very beneficial for language modeling, since speech can be truly regarded as another “natural language” then. However, there is usually a noticeable degradation in reconstruction performance in these tokens. A critical problem lies in how to better utilize the highly-compact discrete VQ representation space. For $F$, the frame rates of most tokens are still much greater than text sequences, which can significantly influence the syntactic and semantic modeling capability of language models [250]. However, reducing the frame rate of tokens may also lead to decreased intelligibility in reconstructed speech. A lower $V$ is also desirable for language modeling and length reduction by BPE.
For acoustic tokens, it remains an open problem what the lower bound of bitrate and the frame rate $F$ are, and how to reach them. More powerful network architectures or advanced VQ strategies should be helpful, and reducing temporal redundancy by disentangling global information is also a promising solution.
VII-2 Streaming Ability and Efficiency
@@
Streaming ability also poses a requirement for model efficiency. Currently, larger acoustic token models are reported to achieve better performance with lower bitrates [78, 81], but at a cost of efficiency. In addition to reducing the bitrate of the tokenized codes, the efficiency of tokenizers must also be balanced for real-time applications.
VII-3 Disentanglement in Acoustic Tokens
@@
VII-4 Variable Frame Rate Tokens
@@
Current speech tokens are usually designed at a fixed temporal rate, while the underlying linguistic units in speech are highly variable-rate. This can offer an important insight and potential of further reducing the bitrate, and more importantly, closing the gap between speech tokens and natural language units for downstream tasks. So far, only semantic tokens have explored varying temporal rates, but with a complicated algorithm or iterative training process [45, 168]. More explorations need to be taken on variable frame rate acoustic tokens and the benefit of these variable frame rate tokens in practice.
VII-5 Combining Acoustic and Semantic Tokens
@@
Given the distinct properties of acoustic and semantic tokens, a natural question arises: Can a representation space contain rich speech understanding capabilities while also reconstructing acoustic details at a decent level? It is promising to incorporating semantic information from SSL models to enhance the reconstruction and downstream modeling performance of acoustic tokens [9, 26, 18]. This approach should be further optimized to achieve lower bitrates and even a single codebook.
VII-6 Paralinguistics in Semantic Tokens
@@
While speaker information is generally considered irrelevant for semantic content, prosody serves as a crucial component of paralinguistic information. Semantic tokens derived through simple clustering methods are likely to discard both speaker information and prosody, harming downstream models’ ability to handle rich emotions, tones, singing voices, and non-verbal vocalizations that convey semantic meaning. This problem can be partially mitigated by certain VQ approaches that encode more information from SSL features [135, 133, 134], but at a cost of more codebooks and higher bitrates. Supervised tokenization could also be considered for directly guiding tokens toward paralinguistic information. However, current supervised tokens rely solely on ASR, which is insufficient for preserving prosody.
VII-7 Noise Preservation vs. Noise Robustness
@@
Similar to disentanglement in acoustic tokens, the inclusion or exclusion of background noise and channel effects in the tokens also depends on the specific application. Most acoustic tokens are designed to capture noise, but their performance across various types of noise and channel effects remains unclear. This issue extends beyond speech and relates to the broader scope of neural audio codecs. On the other hand, denoising [62] is also an interesting application of acoustic tokens that leverages the limited VQ space. If noise is considered undesirable in tokens, such as semantic tokens, then the robustness against various types of signal perturbations needs to be investigated.
VII-8 Timbre Control in Token Vocoders
@@
For semantic tokens and speaker-decoupled acoustic tokens, token vocoders should be responsible for controlling speaker timbre. Currently, both GAN-based token-to-wav vocoders [171] and flow matching-based token-to-mel models [41] have demonstrated strong timbre control capabilities. It remains an open question whether the upper bound of the former method can be improved by training on large-scale datasets, as is done with the latter. Also, the timbre controllability of in-the-wild reference prompts with various acoustic conditions should be further investigated.
VII-9 Adaptivity
The diverse application scenarios range from transmission to various speech processing tasks and speech LLMs. It is challenging to develop a single speech tokenizer that meets all the requirements of these scenarios. However, adaptivity remains a desirable property, where the tokenizer is expected to adapt to different scenarios given a task prompt or a few data examples. If such tokens exist and the aforementioned challenges are well addressed, then we may be close to achieving a universal speech tokenizer.
VIII Conclusion
@@
Recently, discrete speech tokens have emerged as a rapidly evolving field and a core research direction in the speech LLM era. These tokens encode acoustic or semantic information into a compact discrete representation space, catalyzing the fusion of LLMs and speech processing. Existing discrete speech tokens show rich diversity in model architecture and optimization objectives. In this review, we provide a comprehensive introduction to representative categories of discrete speech tokens, summarizing their motivations and limitations. We conduct a unified analysis of reconstruction and voice conversion across different token types to highlight their unique characteristics. We also review efforts to apply discrete tokens to speech processing tasks, including spoken language understanding, speech generation, and spoken language modeling. Finally, we explore future directions for discrete speech tokenization methods. We hope this review lays a solid foundation for future research in speech technology.

### v3

@@ VII-D3 Limitations
VIII Challenges and Future Directions
@@
VIII-1 Low-Bitrate Tokens
For bitrates of tokens, factors $Q$ (number of quantizers) and $F$ (frame rate) play a more important role than $V$ (vocabulary size). Using only a single codebook is very beneficial for language modeling and generation since it frees the need for additional designs for multi-codebook tokens. A critical problem lies in how to better utilize the highly-compact discrete VQ representation space. For $F$, the frame rates of most tokens are still much greater than text sequences, which can significantly influence the syntactic and semantic modeling capability of language models [wang2024whyspeech]. However, there is usually noticeable performance and intelligibility degradation for tokens with single codebook and small $F$. A lower $V$ is also desirable for language modeling and length reduction by BPE.
It remains an open problem what the lower bound of bitrate and the frame rate $F$ are, and how to reach them. More powerful network architectures or advanced VQ strategies should be helpful, and reducing temporal redundancy by disentangling global information is also a promising solution.
VIII-2 Streaming Ability and Efficiency
@@
Streaming ability also poses a requirement for model efficiency. Currently, larger acoustic token models are reported to achieve better performance with lower bitrates [xin2024bigcodec, parker2024scalingtransformerslowbitratehighquality], but at a cost of efficiency. In addition to reducing the bitrate of the tokenized codes, the efficiency of tokenizers must also be balanced for real-time applications.
VIII-3 Disentanglement in Acoustic Tokens
@@
VIII-4 Variable Frame Rate Tokens
@@
As mentioned in Section V-B, the variable-rate nature of linguistic units can offer an important insight for further reducing the bitrate of tokens, and more importantly, closing the gap between speech tokens and natural language units for downstream tasks. More explorations need to be taken on variable frame rate acoustic tokens and the benefit of these variable frame rate tokens in practice.
VIII-5 Combining Acoustic and Semantic Tokens
@@
Given the distinct properties of acoustic and semantic tokens, a natural question arises: Can a representation space contain rich speech understanding capabilities while also reconstructing acoustic details at a decent level? Incorporating semantic information from SSL models has proven to enhance the reconstruction and downstream modeling performance of acoustic tokens [zhang2024speechtokenizer, ye2024codec, liu2024semanticodec]. Recently, explicit text supervision has also sparked remarkable progress in acoustic tokens [qiang2025secousticodec, wang2025tadicodec, gong2025xy]. We anticipate more promising results in this direction.
VIII-6 Paralinguistics in Semantic Tokens
@@
While speaker information is generally considered irrelevant for semantic content, prosody serves as a crucial component of paralinguistic information. Semantic tokens derived through simple clustering methods are likely to discard both speaker information and prosody, harming downstream models’ ability to handle rich emotions, tones, singing voices, and non-verbal vocalizations that convey semantic meaning. This problem can be partially mitigated by certain VQ approaches that encode more information from SSL features [huang2023repcodec, shi24h_interspeech, mousavi2024should], but at a cost of more codebooks and higher bitrates. Supervised tokenization could also be considered for directly guiding tokens toward paralinguistic information in the future.
VIII-7 Noise Preservation vs. Noise Robustness
@@
Similar to disentanglement in acoustic tokens, the inclusion or exclusion of background noise and channel effects in the tokens also depends on the specific application. Most acoustic tokens are designed to capture noise, but their performance across various types of noise and channel effects remains unclear. This issue extends beyond speech and relates to the broader scope of neural audio codecs. On the other hand, denoising [zeghidour2021soundstream] is also an interesting application of acoustic tokens that leverages the limited VQ space. If noise is considered undesirable in tokens, such as semantic tokens, then the robustness against various types of signal perturbations needs to be investigated.
VIII-8 Timbre Control in Token Vocoders
@@
For semantic tokens and speaker-decoupled acoustic tokens, token vocoders should be responsible for controlling speaker timbre. Currently, both GAN-based token-to-wav vocoders [guo2024vec2wav] and flow matching-based token-to-mel models [du2024cosyvoice] have demonstrated strong timbre control capabilities. It remains an open question whether the upper bound of the former method can be improved by training on large-scale datasets, as is done with the latter. Also, the timbre controllability of in-the-wild reference prompts with various acoustic conditions should be further investigated.
IX Conclusion
@@
Recently, discrete speech tokens have emerged as a rapidly evolving field and a core research direction in the speech LLM era. These tokens encode acoustic or semantic information into a compact discrete representation space, catalyzing the fusion of LLMs and speech processing. In this review, we provide a comprehensive introduction to representative categories of discrete speech tokens, summarizing their motivations and limitations. We conduct a unified analysis of reconstruction, voice conversion, and downstream semantic modeling across different token types to highlight their unique characteristics. We also review popular applications of discrete tokens in speech processing tasks, including understanding, generation and language modeling of speech. Finally, we explore future directions for discrete speech tokenization methods. We hope this review lays a solid foundation for future research in speech technology.
@@ Acknowledgments
Appendix I List of Widely-Used Discrete Speech Tokens
To give a high-level overview of existing discrete speech tokens, we summarize the widely-used discrete acoustic tokens in Table II, and semantic tokens in Table III.
TABLE II: A summary of widely-used acoustic speech tokens (neural speech codecs). Note that some codecs in this list are also applicable to general audio. Italic “C,T,U” denote CNN, Transformer or U-Net-based generator architecture. Symbols ‘/’ and ‘-’ denote “or” and “to” for different model versions, “+” means different configurations in different VQ streams in a single model, and $\approx$” means the average value. $Q,F,V$ mean number of quantizers, frame rate and vocabulary size of each quantizer respectively. For example, “$Q=2$, VV=8192+(4096-32768)” in SemantiCodec means one of the two VQ streams has 8192 possible codes, and the other can vary from 4096 to 32768 in different configurations. For FSQ, $V$ is presented as $L^{d}$ where $L$ is the quantization levels for each dimension, and $d$ is the number of dimensions. When $L$ is different for each dimension, like LFSC, we use $L_{1}\times L_{2}\times\cdots$ to represent each dimension. “Dynamic” means the number for each token frame is variable, like in variable-bitrate and variable frame rate setups. Bitrates are computed by $\frac{1}{1000}\sum_{i=1}^{Q}F_{i}\lceil\log_{2}V_{i}\rceil$ kbps, without entropy coding.
General-purpose acoustic tokens
SoundStream [zeghidour2021soundstream] VQ-GAN (C) 24 RVQ max 24 75 1024 max 18.00
EnCodec [encodec] VQ-GAN (C) 24 RVQ max 32 75 1024 max 24.00
TF-Codec [jiang2023latent] VQ-GAN (C) 16 GVQ 3-32 25 512 / 1024 0.68-8.00
Disen-TF-Codec [jiang2023disentangled] VQ-GAN (C) 16 GVQ 2 / 6 25 256 / 1024 0.40 / 1.50
AudioDec [audiodec] VQ-VAE (C)+GAN 48 RVQ 8 160 1024 12.80
HiFi-Codec[yang2023hifi] VQ-GAN (C) 16 / 24 GRVQ 4 50-100 1024 2.00-4.00
DAC [kumar2024high] VQ-GAN (C) 44.1 RVQ 9 86 1024 7.74
LaDiffCodec [yang2024generative] Diffusion 16 RVQ 3 / 6 50 1024 1.50 / 3.00
FreqCodec [du2024funcodec] VQ-GAN (C) 16 RVQ max 32 50 1024 max 16.00
TiCodec [ticodec] VQ-GAN (C) 24 RVQ, GVQ 1-4 75 1024 0.75-3.00
APCodec [APCodec] VQ-GAN (C) 48 RVQ 4 150 1024 6.00
SRCodec [zheng2024srcodec] VQ-GAN (C) 16 GRVQ 2-8 50 512+1024 0.95-3.80
SQ-Codec [yang24l_interspeech] VQ-GAN (C) 16 FSQ 1 50 $19^{32}$ 8.00
Single-Codec [singlecodec] VQ-GAN (T+C) 24 VQ 1 23 8192 0.30
ESC [gu2024esc] VQ-GAN (U) 16 GVQ max 18 50 1024 max 9.00
CoFi-Codec [guo2024speaking] VQ-GAN (U) 16 GVQ 3 8.33+25+50 16384 1.17
HILCodec [ahn2024hilcodec] VQ-GAN (C) 24 RVQ 2-12 75 1024 1.50-9.00
SuperCodec [zheng2024supercodec] VQ-GAN (C) 16 RVQ 2-12 50 1024 1.00-6.00
SNAC [Siuzdak_SNAC_Multi-Scale_Neural_2024] VQ-GAN (C) 24 RVQ 3 12+23+47 4096 0.98
WavTokenizer [ji2024wavtokenizer] VQ-GAN (C) 24 VQ 1 40 / 75 4096 0.48 / 0.90
BigCodec [xin2024bigcodec] VQ-GAN (C) 16 VQ 1 80 8192 1.04
LFSC [casanova2024low] VQ-GAN (C) 22.05 FSQ 8 21.5 $8\times 7\times 6\times 6$ 1.89
NDVQ [niu2024ndvq] VQ-GAN (C) 24 RVQ max 32 75 1024 max 24.00
VRVQ [chae2024variable] VQ-GAN (C) 44.1 RVQ dynamic, max. 8 86 1024 0.26 + max 6.89
TS3-Codec [wu2024ts3codectransformerbasedsimplestreaming] VQ-GAN (T) 16 VQ 1 40 / 50 65536 / 131072 0.64-0.85
Stable-Codec [parker2024scalingtransformerslowbitratehighquality] VQ-GAN (T) 16 FSQ 1 / 2 25 $5^{6}$ / $6^{6}$ 0.40 / 0.70
FreeCodec [zheng2024freecodecdisentangledneuralspeech] VQ-GAN (C+T) 16 VQ 1+1 50+7 256 0.45
FocalCodec [della2025focalcodec] VQ-GAN (T+C) 16 FSQ 1 12.5 / 25 / 50 $2^{13}$ 0.16 / 0.33 / 0.65
ALMTokenizer [yang2025almtokenizer] VQ-GAN (T) 24 RVQ 3 12.5 2048 0.41
TFC [zhang25k_interspeech_tfc] VQ-GAN (C) 24 RVQ 8 dynamic, 18.75-75 1024 1.50-6.00
CodecSlime [wang2025codecslime] VQ-GAN (C) 16 FSQ 1 dynamic, $\approx$40 $5^{2}\times 3^{6}$ 0.08 + ($\approx$0.60)
Mixed-objective acoustic tokens: semantic distillation
Siahkoohi et al. [siahkoohi22_interspeech] VQ-GAN (C) 16 RVQ 2+1 / 2+2 / 6 25+50 64 0.60 / 0.90 / 1.80
SpeechTokenizer [zhang2024speechtokenizer] VQ-GAN (C) 16 RVQ 8 50 1024 4.00
SemantiCodec [liu2024semanticodec] Diffusion 16 VQ 2 12.5-50 8192+(4096-32768) 0.31-1.40
LLM-Codec [yang2024uniaudio15] VQ-GAN (C) 16 RVQ 3 8.33+16.67+33.33 3248+32000+32000 0.85
X-Codec [ye2024codec] VQ-GAN (C) 16 RVQ max 8 50 1024 max 4.00
SoCodec [guo2024socodec] VQ-GAN (C) 16 GVQ 1 / 4 / 8 25 / 8.3 / 4.2 16384 0.35 / 0.47
Mimi [kyutai2024moshi] VQ-GAN (C+T) 24 RVQ 8 12.5 2048 1.10
X-Codec 2.0 [ye2025llasa] VQ-GAN (C+T) 16 FSQ 1 50 $4^{8}$ 0.80
BiCodec [wang2025sparktts] VQ-GAN (C) 16 VQ 1 50 8192 0.65
DualCodec [li25e_dualcodec] VQ-GAN (C) 24 RVQ 3 / 6 12.5 / 25 (1024 / 16384)+(1024 / 4096) 0.75-0.93
XY-Tokenizer [gong2025xy] VQ-GAN (T) 16 RVQ 8 12.5 1024 1.00
TaDiCodec [wang2025tadicodec] Diffusion 24 FSQ 1 6.25 $2^{14}$ 0.0875
Mixed-objective acoustic tokens: disentanglement
SSVC [SSVC] VQ-GAN (C) 24 RVQ 4 50 512 1.80
PromptCodec [pan2024promptcodec] VQ-GAN (C) 24 GRVQ 1-4 75 1024 0.75-3.00
FACodec [facodec] VQ-GAN (C) 16 RVQ 1+2+3 80 1024 4.80
LSCodec [guo2024lscodec] VQ-VAE (C+T)+GAN 24 VQ 1 25 / 50 1024 / 300 0.25 / 0.45
SD-Codec [bie2024learning] VQ-GAN (C) 16 RVQ 12 50 1024 6.00
DeCodec [luo2025decodec] VQ-GAN (C) 16 RVQ 8+8 50 1024 8.00
TABLE III: A high-level summary of widely-used semantic speech tokens. Notations follow Table.II. Symbol ‘/’ denotes different versions. “Inner Quantizer” refers to whether the model has a quantizer, or external quantization (e.g. clustering) must be performed. $F$ denotes frame rate. In case there are inner quantizers, $Q,V$ denote number of quantizers and vocabulary size for each quantizer, respectively. “NR.” means not reported.
From self-supervised learning (SSL) models
vq-wav2vec [vq-wav2vec] Contrastive 0.96k 100 GVQ, $Q=2,V=320$
wav2vec 2.0 [baevski2020wav2vec] Contrastive 60k 50 GVQ, $Q=2,V=320$
XLSR-53 [conneau21_interspeech] Contrastive 50k 50 GVQ, $Q=2,V=320$
HuBERT [hsu2021hubert] Predictive 60k 50 No
WavLM [chen2022wavlm] Predictive 94k 50 No
BEST-RQ [chiu2022self] Predictive 60k 25 No
w2v-BERT [chung2021w2v] Predictive+Contrastive 60k 50 VQ, $Q=1,V=1024$
w2v-BERT 2.0 [barrault2023seamless] Predictive+Contrastive 4500k 50 GVQ, $Q=2,V=320$
DinoSR [liu2024dinosr] Predictive 0.96k 50 VQ, $Q=8,V=256$
NEST-RQ [han2024nest] Predictive 300k 25 No
LAST [turetzky2024last] Predictive NR. 50 VQ, $Q=1,V=500$
Gat et al. [gat2023augmentation] Noise Invariance 0.10k 50 VQ, G=1,V=50G=1,V=50-$500$
ContentVec [qian2022contentvec] Speaker Invariance 0.96k 50 No
SPIRAL [huang2022spiral] Noise Invariance 60k 12.5Hz No
CCC-wav2vec 2.0 [ccc-wav2vec2.0] Noise Invariance 0.36k 50 GVQ, $G=2,V=320$
Spin [chang23_interspeech] Speaker Invariance 0.10k 50 VQ, Q=1,V=128Q=1,V=128-$2048$
NAST [messica2024nast] Noise Invariance 0.96k 50 VQ, Q=1,V=50Q=1,V=50-$200$
DC-Spin [chang2024dc] Speaker Invariance 0.96k 50 VQ, Q=2,V=(50Q=2,V=(50-500)500)+$4096$
Supervised semantic tokens
$\mathcal{S}^{3}$ Tokenizer [du2024cosyvoice] Supervised ASR 172k 25 / 50 VQ, $Q=1,V=4096$
Zeng et al. [zeng2024scaling] Supervised ASR 90k 12.5 VQ, $Q=1,Q=16384$
Du et al. (CosyVoice 2) [cosyvoice2] Supervised ASR 200k 12.5 FSQ, $Q=1,V=3^{8}$
Du et al. (CosyVoice 3)[du2025cosyvoice3] Supervised Multi-Task 530k 25 FSQ, Details NR.
Appendix II Experimental Details
Appendix II-A Vocoder for Reconstruction and Voice Conversion
In the experiments of reconstruction and voice conversion, we train a speech token vocoder for each of the semantic tokens. This speech token vocoder is chosen as CTX-vec2wavα [guo2024lscodec] as the improved version of CTX-vec2wav [du2024unicats]. This vocoder is a timbre-controllable speech token vocoder. The input speech tokens are converted to their corresponding code-vectors first, where code-vectors are concatenated along channel dimension when there are multiple codebooks. Then, these code-vectors are passed to a Conformer [conformer] frontend before a HifiGAN [kong2020hifigan] generator module. The timbre information is provided to the vocoder model from a reference waveform. This reference is passed to a WavLM-Large [chen2022wavlm] model to extract strong timbre-dependent embeddings. These embeddings are extracted from the 6-th layer of this WavLM model, following previous work [knnvc]. Then, they are fed to the Conformer frontend via position-agnostic cross-attention mechanism [du2024unicats].
The training of these CTX-vec2wavα vocoders follows the GAN-based vocoder paradigm where a set of discriminators are employed. We follow the detailed training setup in this repository [^8] https://github.com/cantabile-kwok/vec2wav2.0. The generator of this vocoder consists of $\approx$32M parameters. We train the vocoder for each token up to 1 million steps on 4 GPUs, with a dynamic batch size of $\approx$36s of speech per GPU. The training data is all LibriTTS training splits, which amount to about 585 hours.
After obtaining a vocoder for each semantic token, we can perform reconstruction comparisons on all tokens, and voice conversion comparisons on semantic tokens and some acoustic tokens. For reconstruction, we provide the original utterance as the source of timbre information. One may argue that this setup carries a risk of information leakage, since content information may also be implicitly encoded in timbre representations. An alternative strategy is to use another reference prompt from the same speaker to supply timbre information. However, this essentially becomes a “same-speaker conversion” task rather than reconstruction, and the resulting metrics are not strictly comparable to those of ordinary acoustic tokens that do not require additional speaker inputs. Given that the CTX-vec2wavα vocoder has never been trained on cases where timbre representations and content tokens originate from the same segment, we consider this strategy to still provide sufficiently fair comparisons.
The details in evaluation metrics include:
The test metadata for reconstruction and voice conversion can be found here [^16] https://cpdu.github.io/unicats/resources/testsetB_utt2prompt and here [^17] https://cantabile-kwok.github.io/LSCodec/audio_ready/VC/utt2prompt, respectively.
Appendix II-B Downstream Semantic Modeling
Appendix II-B1 Motivation
Given that a large branch of discrete speech tokens is semantic tokens, it is also important to measure the semantic modeling abilities of speech tokens in downstream tasks, without relying on a codec decoder or vocoder to reconstruct into waveforms. There are mainly two approaches to probe the semantic modeling abilities in discrete speech tokens:
  • Index-based: Like DASB [mousavi2024dasb], discrete indices are directly used as inputs and passed to a set of learnable embedding lookup modules before subsequent networks. When there are multiple code streams (for example, when using GVQ or RVQ), a learnable attention weight will aggregate the embeddings from different code streams.
  • Vector-based: The code-vectors corresponding to the discrete indices are considered as inputs to the probing network. This is a direct extension to the SUPERB [superb] benchmark that mainly considers continuous speech representations.
The vector-based approach, in other words, probes the information from the established latent space where speech tokenizers perform quantization on. Meanwhile, the index-based approach concentrates on the abstract discrete indices and anticipates that indices themselves carry important semantic information rather than the underlying high-dimensional space. Since the most of the discrete token-based speech generation models use token indices instead of code-vectors as modeling targets, we opt to use the index-based approach in this survey.
To better align with the results in previous works, we use the DASB [mousavi2024dasb] framework and its official implementation [^18] https://github.com/speechbrain/benchmarks/tree/DASB/benchmarks/DASB.
Appendix II-B2 Tasks and Datasets
We consider two of the most representative speech semantic tasks: automatic speech recognition (ASR) and intent classification (IC). In ASR, we use the characters as text units for simplicity, and the connectionist temporal classification (CTC) [ctc] criterion for training. We use all the 960 hours in LibriSpeech [librispeech] as the training data, dev-clean split as the validation set, and test-clean as the test set. IC is the process of understanding the underlying goal or purpose of a user’s spoken utterance. We use SLURP [bastianelli-etal-2020-slurp] for the IC task, which includes 18 classification categories, such as event name, date, etc. We use the train-real split for training and the official dev and test sets, which contains about 40, 7 and 10 hours of speech data, respectively.
All the original waveforms are in 16kHz. If a speech tokenizer requires input at a different sampling rate, we resample the waveform before feeding them accordingly.
Appendix II-B3 Model
Following DASB, we use LSTM-based networks for both tasks. For each token, the input layer consists of $Q$ embedding lookup tables of shape $V\times 1024$, where $Q$ is the number of codebooks and $V$ is the vocabulary size of each codebook. For each frame, the $Q$ embeddings are fed into an multi-layer perceptron (MLP) with a hidden layer of 1024 dimensions, to produce softmax probabilities that weight-sum the $Q$ embeddings. Then, an bidirectional LSTM with 2 layers and 1024 hidden neurons takes the embedding sequence as input, and produces 2048-dimensional outputs. For the ASR task, another linear layer is applied to produce emission probabilities on the 31 text units (26 English characters and special tokens). For the IC task, mean pooling over the sequence axis is applied before a linear projection to the logits on 18 output classes. Besides the input embedding layer which is dependent on the specific token configuration, the model has around 43M parameters.
It is important to note that some speech tokens have a low frame rate, such as 12.5Hz for Mimi [kyutai2024moshi]. As the character rate is about 14Hz, the original low frame rates make these tokens not suitable for character-based CTC ASR. Hence, when the token frame rate is less than 50Hz, we always repeat the input tokens to at least 50Hz for ASR task. For the IC task, no repeating is performed.
Appendix II-B4 Training and Decoding
For the ASR task, we use AdamW [loshchilov2018decoupled] optimizer with initial learning rate $10^{-4}$ and weight decay $5\times 10^{-4}$. For the IC task, we use Adam [adam] optimizer with initial learning rate $2\times 10^{-4}$. For both tasks, the learning rate is reduced to 80% of its original value whenever the improvement in validation loss after an epoch is less than 0.25% relative to the previous epoch. For each token, training is conducted for 20 epochs on a single GPU, and the checkpoint with the lowest validation loss is selected for testing.
In ASR task, we use beam search decoding with beam size 100. Other hyperparameters all remain consistent with the DASB official configuration. For some tokens, training suffers from collapse, leading to repetitive and meaningless characters in decoding. This is usually observed when the token has a large vocabulary size (like 2 groups of 15625 FSQ indices for StableCodec [parker2024scalingtransformerslowbitratehighquality], and 2 groups of 32000 VQ indices for LLM-Codec [yang2024uniaudio15]). Following DASB, we denote these cases as “not converged” (NC), since the metrics are not meaningful there.